OpenStudy (anonymous):

What am I doing wrong with this problem? F(x)=x^2+2x y=x^2+2x x^2=y-2x x=y-2 y=x-2 f^-1(x)=x-2 f^-1(3)=3-2

OpenStudy (anonymous):

4th and 5th lines !!!

OpenStudy (anonymous):

after x^2=y - 2x how can u say x=y-2

OpenStudy (anonymous):

oh, that was it. sorry, thanks